OpenStudy (anonymous):

One mole of contain the largest number of atoms A-Al2(SO4)3 B-C10H8 C-Cl2 D-Na3Po4 E-S8

OpenStudy (anonymous):

B-C10H8

OpenStudy (anonymous):

@missiey So the correct answer is A?

OpenStudy (anonymous):

@missiey @rushwr Thanks:)

OpenStudy (whpalmer4):

you have the same number of molecules in each case: about \(6.022*10^{23}\) look at each molecule and count up the atoms to find the molecule that will give you the most atoms in a mole of that molecule. A: \(\text{Al}_2(\text{SO}_4)^3 \rightarrow 2+3*(1+4*1)= 17\) B: \(\text{C}_{10}\text{H}_8\rightarrow 10+8 = 18\) C: \(\text{Cl}_2\rightarrow 2\) D: \(\text{Na}_3\text{PO}_4\rightarrow 3+4=7\) D:\(\text{S}_8\rightarrow8\) Therefore, B is the correct answer, NOT A.